OpenStudy (anonymous):

let (an) and (bn) be two non negative sequences. Let cn=max{an,bn} for each n∈N. Prove the following statements or give a counter example when false (a) If (an) and (bn) converge, so does (cn). (b)If (an)and(bn) diverge, so does(cn c) if the series of an and the series of bn converges then so does the series of cn. d) if the series of an and the series of bn diverges then so does the series of cn

OpenStudy (anonymous):

I'm thinking these are both false?

OpenStudy (anonymous):

Do u know this one @Zarkon ?

OpenStudy (dumbcow):

this is same as last question?

OpenStudy (zarkon):

yes...i do....again what have you tried?

OpenStudy (anonymous):

Its similar .. this is the summation of the an and bn

OpenStudy (anonymous):

@Zarkon well I know that in order for a series to converge it must me bounded above correct?

OpenStudy (zarkon):

the partial sums were bounded above then yes it would converge

OpenStudy (anonymous):

Do we have to use anything with the Cauchy criterion or no?

OpenStudy (zarkon):

you don't need to

OpenStudy (anonymous):

Was I correct in saying these are both false at least?

OpenStudy (zarkon):

though you might want to

OpenStudy (anonymous):

They are both false?

OpenStudy (zarkon):

no they are not both false

OpenStudy (anonymous):

I posted a and b to this question earlier and two people helped me.. the questions were if (an) and (bn) both converge then (cn) converges as well..and the same for diverge.. they told me they were both true..is that in correct?

OpenStudy (zarkon):

we are given that \[\sum_{n=1}^{\infty}a_n\text{ and }\sum_{n=1}^{\infty}b_n\] both converge correct

OpenStudy (zarkon):

for (c)

OpenStudy (anonymous):

Yes correct..without the limits

OpenStudy (anonymous):

I edited the question so u could see a and b as well

OpenStudy (zarkon):

(a) is true...(b) is not true

OpenStudy (anonymous):

So then c is true and d is not true?

OpenStudy (zarkon):

c and d are also true

OpenStudy (anonymous):

ayi yi I don't understand any of this.. I am going to lose internet connect for a little bit. Will u please help m,e when I get nback ..Id really appreciate it

OpenStudy (zarkon):

for (c) and (d) you can use the comparison test

OpenStudy (anonymous):

What about for a? If I just show that can is equal to bn or an and thy both convene then cn also converges?

OpenStudy (zarkon):

you really need two cases...when (1) an->a and bn->b and a=b 2) if a>b or b>a without loss of generality you can assume a>b

OpenStudy (anonymous):

Ok and how would I go about disproving b?

OpenStudy (zarkon):

counter example: \[a_n=\left\{\begin{array}{ll} 1, & n \text{ even} \\ 2, & n \text{ odd}\\ \end{array}\right.\] \[b_n=\left\{\begin{array}{ll} 2, & n \text{ even} \\ 1, & n \text{ odd}\\ \end{array}\right.\] \[c_n=2\]

OpenStudy (zarkon):

do you understand?

OpenStudy (anonymous):

Umm not exactly

OpenStudy (zarkon):

\[\lim_{n\to \infty}a_n\] does not converge correct (ie it diverges)

OpenStudy (zarkon):

using my \(a_n\)

OpenStudy (zarkon):

the sequence (starting at n=1) is 2,1,2,1,2,1,2,1,2,1,2,1..... it does not get close to one number

OpenStudy (zarkon):

same with \(b_n\) 1,2,1,2,1,2,1,2,1,2,1,2,1,2...

OpenStudy (zarkon):

but \(c_n\) is the max of the two sequences at any value n...that is 2 so \[\lim_{n\to \infty}c_n=2\]

OpenStudy (anonymous):

So therefor it doesn't converge ?

OpenStudy (zarkon):

\(\{a_n\}\) and \(\{b_n\}\) both diverge but \(\{c_n\}\) converges

OpenStudy (zarkon):

here we are talking about sequences not series

OpenStudy (anonymous):

ok got i get that part

OpenStudy (anonymous):

and for part c i use the comparisson test that says if series bk converges then ak converges and since cn is the max then it must converge as well?

OpenStudy (anonymous):

i only know the proof of comparisson test for seqences not series..is it alot different?

OpenStudy (zarkon):

if \[\sum_{n=1}^{\infty}a_n\text{ and }\sum_{n=1}^{\infty}b_n\] converge then \[\sum_{n=1}^{\infty}[a_n+b_n]\] converges and since \[c_n=max\{a_n,b_n\}\le a_n+b_n\] we have that \[\sum_{n=1}^{\infty}c_n\] converges by the comparison test

OpenStudy (anonymous):

ahh thank you!

OpenStudy (anonymous):

@zarkon so then d is exactly the same just with divergence?

OpenStudy (zarkon):

if \[\sum_{n=1}^{\infty}a_n\] diverges and \(a_n\le c_n\) then \[\sum_{n=1}^{\infty}c_n\] diverges by the comparison test